# KiDS-SBI: Simulation-Based Inference Analysis of KiDS-1000 Cosmic Shear

### v1

@@ Acknowledgements.
We thank Stephen Feeney, Justin Alsing, Hendrik Hildebrandt, Marika Asgari and Alan Heavens for helpful discussions. MvWK thanks the Science and Technology Facilities Council for support. KL thanks the UCL Cosmoparticle Initiative. NT and BJ are supported by the ERC-selected UKRI Frontier Research Grant EP/Y03015X/1 and the UK Space Agency grants ST/W002574/1 and ST/X00208X/1. BJ is also supported by STFC Consolidated Grant ST/V000780/1. AL acknowledges support from the research project grant ‘Understanding the Dynamic Universe’ funded by the Knut and Alice Wallenberg Foundation under Dnr KAW 2018.0067. RR is supported by the European Research Council (Grant No. 770935). This work used facilities provided by the UCL Cosmoparticle Initiative. We thank Dr Edd Edmonson for technical support at the UCL HPC facilities. The results in this paper are based on observations made with ESO Telescopes at the La Silla Paranal Observatory under programme IDs: 088.D-4013, 092.A-0176, 092.D-0370, 094.D-0417, 177.A-3016, 177.A-3017, 177.A-3018 and 179.A-2004, and on data products produced by the KiDS consortium. The KiDS production team acknowledges support from: Deutsche Forschungsgemeinschaft, ERC, NOVA and NWO-M grants; Target; the University of Padova, and the University Federico II (Naples). Contributions to the data processing for VIKING were made by the VISTA Data Flow System at CASU, Cambridge and WFAU, Edinburgh. The work presented in this paper made use of the following software packages: Numpy (Harris et al., 2020), SciPy (Virtanen et al., 2020), AstroPy (Astropy Collaboration et al., 2018), HealPy (Zonca et al., 2019), GetDist (Lewis, 2019), Jupyter Lab (Kluyver et al., 2016), numba (Lam et al., 2015), and Matplotlib (Hunter, 2007).
@@ Appendix A Numerical non-Limber integration
where $i,q=1,...,d;j,m=1,...,n$ Any basis is suitable, here we chose equidistant collocation points $x_{j}$ and the $n$ lowest-order polynomials as basis functions. Finally, the vector ${\bf w}$ depends on the specific type of integral considered. Eq. 8 allows for two possibilities: $(i)$ first integrate over the co-moving distance $\chi$: $(ii)$ first integrate over the wave-vector $k$. Both approaches have advantages and disadvantages: due to the separation of the $\chi$ and $\chi^{\prime}$ dependence, integrating over the co-moving distance first will be quicker. However, the residual integral over $k$ can still be very oscillatory and must be sampled accordingly, especially with narrow window functions $W(\chi)$. The second option is to integrate over $k$ first, this smooths out the oscillations from the Bessel function. However, there are two remaining integrals, thus slowing down the computation, especially for many radial bins. If, however, the radial bins do not overlap, the resulting spectrum for bins further apart than a single bin index will be very small and can be neglected. A more detailed discussion on this is provided in Leonard et al. (2023). For integrals of the type
@@ Appendix B Accuracy of the cosmic shear signal
To test the robustness of the cosmic shear signal as predicted by the log-normal simulations generated in KiDS-SBI in addition to the self-consistency tests already conducted within Sect. 3, we perform a comparison of the noise-free pseudo-Cls as measured in KiDS-SBI assuming the standard isotropic systematics model (see Sect. 4.2 and 7.2) to the pseudo-Cls predicted from analytical theory (see Appendix D). This comparison is shown for the KiDS-SBI MAP cosmology in Fig. 21. When assuming 19 shells and a comoving-volume weighted kernel as discussed in Sect. 3.2, we find per-cent or sub-percent level agreement for all tomographic bin combinations and scales (within $0.5\sigma$ or less). The agreement between the models is also shown to be good in Fig. 17 where the posterior from the SBI analysis based on KiDS-SBI is unbiased with respect to the posterior sampled with an MCMC while assuming analytical theory. Overall, we find that for small $\ell$ the forward simulations predict additional power with respect to the pseudo-Cls predicted from theory. This can be explained by discretisation effects along the line-of-sight due to the division of the observable Universe into concentric shells weighted by a window function (see Sect. 3.2), and by the linear interpolation of convergence and shear across HEALPix pixels which are built into SALMO. Both effects have been shown to produce a similar minor excess in angular power in testing conducted in Tessore et al. (2023).
When increasing the resolution along the line-of-sight by increasing the number of shells, and when changing the kernel to a uniform window function as suggested in Tessore et al. (2023), we find that for 26 shells the agreement with the signal predicted by analytical theory is marginally better, but not enough to impact the SBI. When increasing the number of shells further to 31 shells, such that each shell has a width of ${\sim}100$ Mpc $h^{-1}$, the agreement with analytical theory is poorer for large $\ell$ and for low $\ell$ in the autocorrelations, as discretisation effects become important. Consequently, we conclude that the choice of 19 shells allows to accurately forward model the cosmic shear signal in KiDS-SBI, while still allowing for computational efficiency (see Sect. 5.1).
These effects are not significant enough to bias the cosmological inference at the precision of KiDS-1000, particularly, since they are more prevalent in tomographic bin combinations with a lower signal-to-noise ratio. However, future simulation-based inference analyses of galaxy surveys with smaller uncertainties (e.g. Stage-IV surveys) based on similar simulations should aim to account for such effects.
@@ Appendix C Effects of variable depth and shear biases
Figure 22: The impact of variable depth in KiDS-1000 on the cosmic shear real-space two-point correlation functions, $\xi(\theta)$ in the autocorrelation of the S5 bin. The blue line shows the difference in the mean signal from 5,000 realisations of the anisotropic systematics model with variable depth versus the mean signal from the same number of realisations when assuming that the observational depth is isotropic. The orange line shows the same difference in signal as predicted for KiDS-1000 by the analytical model in Heydenreich et al. (2020). Note that variable depth constitutes approximately $1\%$ of the measured cosmic shear signal in KiDS-1000.
Additionally, when modelling variable depth, the shape noise becomes anisotropic across the sky as a function of the variations in the effective galaxy density and the intrinsic galaxy shape dispersion. When considering a Gaussian likelihood, this implies that the shape noise scales with $\sigma_{\epsilon}^{2}/n_{\mathrm{eff}}$ (Kaiser 1992, 1998). As discussed in Sect. 3.8 both quantities scale linearly with the variable depth measure, $\sigma_{\mathrm{rms}}$, which implies that the shape noise in a given tomographic bin either scales with $\sigma_{\mathrm{rms}}^{3}$ in S1, S2 and S5, or with $\sigma_{\mathrm{rms}}^{-1}$ in S3 and S5. Since in KiDS-1000 the distribution of $\sigma_{\mathrm{rms}}$ is skewed below the mean (see Fig. 23), variable depth in fact reduces the overall contribution of the shape noise to the uncertainty of the cosmic shear signal. Nevertheless, as shown in Figs. 7 and 8, variable depth also causes additional non-cosmological angular correlations in the data at the scales of pixels, pointing overlaps, whole pointings or even over the whole footprint, while also changing the signal along the line-of-sight anisotropically as shown in Fig. 10. These additional correlations can significantly alter the covariance of the two-point statistic or may even lead to non-Gaussian noise. Such correlations can exacerbate the uncertainty contributions from cosmic variance and super-sample covariance, which causes the noise predicted in KiDS-1000 by a model including variable depth to ultimately exceed the noise predicted by a model which does not consider the systematic (Reischke et al. in prep.)).
@@ Appendix D Theoretical signal modelling
where $n_{\mathrm{S}}^{(p)}$ is the redshift distribution of the objects in $p^{th}$ tomographic bin as shown in Figure 2. For consistency with the forward modelling described in Sect. 3.5, we choose an intrinsic alignments kernel in accordance with the non-linear alignment model (NLA; Catelan et al. 2001; Hirata et al. 2004; Bridle & King 2007) given by
@@
The first mixing matrix from the left in Eq. (59), $M_{\ell\ell^{\prime\prime}}$, is the standard block mixing matrix to account for the partial sky coverage in a survey (Hivon et al. 2002; Bennett et al. 2003; Brown et al. 2005). The second mixing matrix, $M^{\prime(pq)}_{\ell^{\prime\prime}\ell^{\prime}}$, accounts for an additional effect: the mode mixing induced by a mask of uncorrelated noise originating from the random sampling described in Sect. 3.6. In other words, when randomly Poisson sampling galaxies within a discrete pixel $k$ given a probability of $P^{(m)}(N)$, there will be a non-zero probability of observing zero counts in that given pixel. In essence, all the pixels which are sampled to have zero galaxy counts impose an additional masking on our sample. This can become significant at the resolutions and galaxy densities which we probe in this work. For example, for an HealPix $N_{\mathrm{side}}=2048$ and a galaxy density of $n=1\,\mathrm{arcmin}^{-2}$ for a given tomographic bin $k$ (similar to KiDS-1000), $P^{(m)}(N=0)\approx 0.07$. This means that nearly a tenth of all pixels could be ”masked” in a given realisation of simulations at a higher resolution. To allow for these cases, we decide to also model the mode mixing induced by the sampling variance. However, for the settings discussed in this paper for the main pipeline of forwards simulations where $N_{\mathrm{side}}=1024$, this effect is not so relevant. Even for the sparsest source bin, S1, with an $n_{\mathrm{gal}}^{(1)}=0.62\,\mathrm{\mathrm{arcmin}}^{-2}$, $P^{(m)}(N=0)\approx 0.0005$, so only a negligible fraction of pixels are masked due to random sample variance.
@@
where $\mu^{(pq)}=P^{(p)}(N>0)P^{(q)}(N>0)$, i.e. the product of the mean probabilities of success within each tomographic bin (the probability that a given pixel will be populated by at least one galaxy), while $2\sigma^{(pq)2}=P^{(p)2}(N>0)(1-P^{(p)}(N>0))^{2}+P^{(q)2}(N>0)(1-P^{(q)}(N>0))^{2}$, i.e. the associated standard deviation of the mean, and $N_{\mathrm{pix}}$ is the number of pixels.
@@
$W_{\ell\ell^{\prime}mm^{\prime}}^{+}=\frac{1}{2}(_{2}W_{\ell\ell^{\prime}}^{mm^{\prime}}+_{-2}W_{\ell\ell^{\prime}}^{mm^{\prime}})\,,$ (66)
$W_{\ell\ell^{\prime}mm^{\prime}}^{-}=\frac{\mathrm{i}}{2}(_{2}W_{\ell\ell^{\prime}}^{mm^{\prime}}-_{-2}W_{\ell\ell^{\prime}}^{mm^{\prime}})\,,$ (67)
@@ Appendix E Parameter recovery
Fig. 26 shows the posterior distribution of the seven main parameters varied in the analysis as learnt from 18,000 realisations of the anisotropic systematics model. The data vector used for this posterior is a noisy mock data vector generated with the anisotropic systematics model while assuming the input parameters given in Table 2. We find that the posteriors from KiDS-SBI accurately recover the true input parameters (see Table 4 for the inferred parameter values). As shown in Fig. 26, the anisotropic systematics posterior mostly only constrains $S_{8}$ and $A_{\mathrm{IA}}$. At the same time, it becomes apparent that the posterior distributions over $\omega_{\mathrm{b}}$, $n_{\mathrm{s}}$, $A_{\mathrm{bary}}$ and $h_{0}$ are mainly prior driven, i.e. the posterior distribution is mostly flat throughout prior space. As the cosmic shear signal is largely degenerate to small changes in these parameters, this is expected.
@@
To determine the origin of the observed cosmology dependence in the likelihood in Fig. 18, we investigate whether there is a theoretical justification for the uncertainty in the cosmic shear signal in KiDS-1000 to vary significantly with $S_{8}$. First, we assume that the likelihood of a full-sky shear angular power spectrum, $C(\ell)$, at a given cosmology, is Gaussian, i.e. $\mathcal{L}\propto\mathrm{exp}(-\frac{1}{2}\chi^{2})$, where $\chi$ is the goodness of fit given by
@@
where $R^{(pq)}(\ell)\equiv C_{\mathrm{fid}}^{(pq)}(\ell)/\sqrt{C_{\mathrm{fid}}^{(pp)}(\ell)\,C_{\mathrm{fid}}^{(qq)}(\ell)}$. By calculating the distribution of $\mathrm{exp}(-\frac{1}{2}\chi_{a}^{2})$ over $\chi$ for different values of $\ell$ and $a$, we can assess the impact of rescaling the shear angular power spectrum directly on the Gaussian likelihood across different angular scales, $\ell$. To facilitate this, one can analytically compute the standard deviation of this distribution, $\sigma_{\chi_{a}}$, as follows
@@
Appendix G Posterior variations with noise realisations
As discussed in Sect. 6, different uncertainty realisations of the anisotropic systematics model in KiDS-SBI at the same cosmology produces a significant random scatter in the measured final pseudo-Cls. Fig. 28 shows how the choice of noise realisation impacts the inferred posteriors at a given cosmology. We find that overall the variations in the posteriors between uncertainty realisations are small with the mean marginal of $S_{8}$ fluctuating by up to ±$\pm 0.5\sigma$ and the $A_{\mathrm{IA}}$ marginal can fluctuate up to ±$\pm 0.75\sigma$.
Figure 28: Posterior contours from five different mock data vectors given the anisotropic systematics model within KiDS-SBI over the prior space shown in Table 2. Each data realisation assumes the same input parameters (as given by the black line and shown in Table 4) while varying the random seeds of the underlying matter fields, galaxy positions, intrinsic galaxy shapes, and shear biases. The posteriors shown in Fig. 26,15, 16, and 17 are based on the 3817 noise realisation.
Appendix H Impact of the sampler on the Gaussian likelihood posteriors

### v2

@@ Acknowledgements.
We thank Stephen Feeney, Justin Alsing, Hendrik Hildebrandt, Marika Asgari and Alan Heavens for helpful discussions. We are grateful to the referee for the constructive comments on the manuscript. MvWK thanks the Science and Technology Facilities Council for support. KL thanks the UCL Cosmoparticle Initiative. NT and BJ are supported by the ERC-selected UKRI Frontier Research Grant EP/Y03015X/1 and the UK Space Agency grants ST/W002574/1 and ST/X00208X/1. BJ is also supported by STFC Consolidated Grant ST/V000780/1. AL acknowledges support from the research project grant ‘Understanding the Dynamic Universe’ funded by the Knut and Alice Wallenberg Foundation under Dnr KAW 2018.0067. RR is supported by the European Research Council (Grant No. 770935). AHW is supported by the Deutsches Zentrum für Luft- und Raumfahrt (DLR), made possible by the Bundesministerium für Wirtschaft und Klimaschutz, and acknowledges funding from the German Science Foundation DFG, via the Collaborative Research Center SFB1491 ’Cosmic Interacting Matters - From Source to Signal’. This work used facilities provided by the UCL Cosmoparticle Initiative. We thank Dr Edd Edmonson for technical support at the UCL HPC facilities. The results in this paper are based on observations made with ESO Telescopes at the La Silla Paranal Observatory under programme IDs: 088.D-4013, 092.A-0176, 092.D-0370, 094.D-0417, 177.A-3016, 177.A-3017, 177.A-3018 and 179.A-2004, and on data products produced by the KiDS consortium. The KiDS production team acknowledges support from: Deutsche Forschungsgemeinschaft, ERC, NOVA and NWO-M grants; Target; the University of Padova, and the University Federico II (Naples). Contributions to the data processing for VIKING were made by the VISTA Data Flow System at CASU, Cambridge and WFAU, Edinburgh. The work presented in this paper made use of the following software packages: Numpy (Harris et al., 2020), SciPy (Virtanen et al., 2020), AstroPy (Astropy Collaboration et al., 2018), HealPy (Zonca et al., 2019), GetDist (Lewis, 2019), Jupyter Lab (Kluyver et al., 2016), numba (Lam et al., 2015), and Matplotlib (Hunter, 2007).
@@ Appendix A Numerical non-Limber integration
where $i,q=1,...,d;j,m=1,...,n$ Any basis is suitable, here we chose equidistant collocation points $x_{j}$ and the $n$ lowest-order polynomials as basis functions. Finally, the vector ${\bf w}$ depends on the specific type of integral considered. Equation 8 allows for two possibilities: $(i)$ first integrate over the co-moving distance $\chi$: $(ii)$ first integrate over the wave-vector $k$. Both approaches have advantages and disadvantages: due to the separation of the $\chi$ and $\chi^{\prime}$ dependence, integrating over the co-moving distance first will be quicker. However, the residual integral over $k$ can still be very oscillatory and must be sampled accordingly, especially with narrow window functions $W(\chi)$. The second option is to integrate over $k$ first, this smooths out the oscillations from the Bessel function. However, there are two remaining integrals, thus slowing down the computation, especially for many radial bins. If, however, the radial bins do not overlap, the resulting spectrum for bins further apart than a single bin index will be very small and can be neglected. A more detailed discussion on this is provided in Leonard et al. (2023). For integrals of the type
@@ Appendix B Accuracy of the cosmic shear signal
To test the robustness of the cosmic shear signal as predicted by the log-normal simulations generated in KiDS-SBI in addition to the self-consistency tests already conducted within Sect. 3, we perform a comparison of the noise-free pseudo-Cls as measured in KiDS-SBI assuming the standard isotropic systematics model (see Sects. 4.2 and 7.2) to the pseudo-Cls predicted from analytical theory (see Appendix D). This comparison is shown for the KiDS-SBI MAP cosmology in Fig. 20. When assuming 19 shells and a comoving-volume weighted kernel as discussed in Sect. 3.2, we find per-cent or sub-percent level agreement for all tomographic bin combinations and scales (within $0.5\sigma$ or less). The agreement between the models is also shown to be good in Fig. 16 where the posterior from the SBI analysis based on KiDS-SBI is unbiased with respect to the posterior sampled with an MCMC while assuming analytical theory. Overall, we find that for small $\ell$ the forward simulations predict additional power with respect to the pseudo-Cls predicted from theory. This can be explained by discretisation effects along the line of sight due to the division of the observable Universe into concentric shells weighted by a window function (see Sect. 3.2), and by the linear interpolation of convergence and shear across HEALPix pixels which are built into SALMO. Both effects have been shown to produce a similar minor excess in angular power in testing conducted in Tessore et al. (2023).
When increasing the resolution along the line of sight by increasing the number of shells, and when changing the kernel to a uniform window function as suggested in Tessore et al. (2023), we find that for 26 shells the agreement with the signal predicted by analytical theory is marginally better, but not enough to impact the SBI. When increasing the number of shells further to 31 shells, such that each shell has a width of ${\sim}100$ Mpc $h^{-1}$, the agreement with analytical theory is poorer for large $\ell$ and for low $\ell$ in the autocorrelations, as discretisation effects become important. Consequently, we conclude that the choice of 19 shells allows us to carry out an accurate forward modelling of the cosmic shear signal in KiDS-SBI, while still allowing for computational efficiency (see Sect. 5.1).
These effects are not significant enough to bias the cosmological inference at the precision of KiDS-1000, particularly, since they are more prevalent in tomographic bin combinations with a lower signal-to-noise ratio. However, future simulation-based inference analyses of galaxy surveys with smaller uncertainties (for instance stage-IV surveys) based on similar simulations should aim to account for such effects.
@@ Appendix C Effects of variable depth and shear biases
Figure 21: The impact of variable depth in KiDS-1000 on the cosmic shear real-space two-point correlation functions, $\xi(\theta)$ in the autocorrelation of the S5 bin. The blue line shows the difference in the mean signal from 5,000 realisations of the anisotropic systematics model with variable depth versus the mean signal from the same number of realisations when assuming that the observational depth is isotropic. The orange line shows the same difference in signal as predicted for KiDS-1000 by the analytical model in Heydenreich et al. (2020). We note that variable depth constitutes approximately $1\%$ of the measured cosmic shear signal in KiDS-1000.
Additionally, when modelling variable depth, the shape noise becomes anisotropic across the sky as a function of the variations in the effective galaxy density and the intrinsic galaxy shape dispersion. When considering a Gaussian likelihood, this implies that the shape noise scales with $\sigma_{\epsilon}^{2}/n_{\mathrm{eff}}$ (Kaiser 1992, 1998). As discussed in Sect. 3.8 both quantities scale linearly with the variable depth measure, $\sigma_{\mathrm{rms}}$, which implies that the shape noise in a given tomographic bin either scales with $\sigma_{\mathrm{rms}}^{3}$ in S1, S2 and S5, or with $\sigma_{\mathrm{rms}}^{-1}$ in S3 and S5. Since in KiDS-1000 the distribution of $\sigma_{\mathrm{rms}}$ is skewed below the mean (see Fig. 22), variable depth in fact reduces the overall contribution of the shape noise to the uncertainty of the cosmic shear signal. Nevertheless, as shown in Figs. 7 and 8, variable depth also causes additional non-cosmological angular correlations in the data at the scales of pixels, pointing overlaps, whole pointings or even over the whole footprint, while also changing the signal along the line of sight anisotropically as shown in Fig. 10. These additional correlations can significantly alter the covariance of the two-point statistic or may even lead to non-Gaussian noise. Such correlations can exacerbate the uncertainty contributions from cosmic variance and super-sample covariance, which causes the noise predicted in KiDS-1000 by a model including variable depth to ultimately exceed the noise predicted by a model which does not consider the systematic (Reischke et al. 2024)).
@@ Appendix D Theoretical signal modelling
where $n_{\mathrm{S}}^{(p)}$ is the redshift distribution of the objects in $p^{th}$ tomographic bin as shown in Fig. 2. For consistency with the forward modelling described in Sect. 3.5, we chose an intrinsic alignment kernel in accordance with the non-linear alignment model (NLA; Catelan et al. 2001; Hirata et al. 2004; Bridle & King 2007) given by
@@
The first mixing matrix from the left in Eq. (59), $M_{\ell\ell^{\prime\prime}}$, is the standard block mixing matrix to account for the partial sky coverage in a survey (Hivon et al. 2002; Bennett et al. 2003; Brown et al. 2005). The second mixing matrix, $M^{\prime(pq)}_{\ell^{\prime\prime}\ell^{\prime}}$, accounts for an additional effect: the mode mixing induced by a mask of uncorrelated noise originating from the random sampling described in Sect. 3.6. In other words, when randomly Poisson sampling galaxies within a discrete pixel $k$ given a probability of $P^{(m)}(N)$, there will be a non-zero probability of observing zero counts in that given pixel. In essence, all the pixels which are sampled to have zero galaxy counts impose an additional masking on our sample. This can become significant at the resolutions and galaxy densities which we probe in this work. For example, for an HealPix $N_{\mathrm{side}}=2048$ and a galaxy density of $n=1\,\mathrm{arcmin}^{-2}$ for a given tomographic bin $k$ (similar to KiDS-1000), $P^{(m)}(N=0)\approx 0.07$. This means that nearly a tenth of all pixels could be ’masked’ in a given realisation of simulations at a higher resolution. To allow for these cases, we decide to also model the mode mixing induced by the sampling variance. However, for the settings discussed in this paper for the main pipeline of forwards simulations where $N_{\mathrm{side}}=1024$, this effect is not so relevant. Even for the sparsest source bin, S1, with an $n_{\mathrm{gal}}^{(1)}=0.62\,\mathrm{\mathrm{arcmin}}^{-2}$, $P^{(m)}(N=0)\approx 0.0005$, so only a negligible fraction of pixels are masked due to random sample variance.
@@
where $\mu^{(pq)}=P^{(p)}(N>0)P^{(q)}(N>0)$, namely the product of the mean probabilities of success within each tomographic bin (the probability that a given pixel will be populated by at least one galaxy), while $2\sigma^{(pq)2}=P^{(p)2}(N>0)(1-P^{(p)}(N>0))^{2}+P^{(q)2}(N>0)(1-P^{(q)}(N>0))^{2}$, expressly the associated standard deviation of the mean, and $N_{\mathrm{pix}}$ is the number of pixels.
@@
$\begin{split}W_{\ell\ell^{\prime}mm^{\prime}}^{+}=\frac{1}{2}(_{2}W_{\ell\ell^{\prime}}^{mm^{\prime}}+_{-2}W_{\ell\ell^{\prime}}^{mm^{\prime}})\,,\\ W_{\ell\ell^{\prime}mm^{\prime}}^{-}=\frac{\mathrm{i}}{2}(_{2}W_{\ell\ell^{\prime}}^{mm^{\prime}}-_{-2}W_{\ell\ell^{\prime}}^{mm^{\prime}})\,,\end{split}$ (66)
@@ Appendix E Parameter recovery
Figure 25 shows the posterior distribution of the seven main parameters varied in the analysis as learnt from 18,000 realisations of the anisotropic systematics model. The data vector used for this posterior is a noisy mock data vector generated with the anisotropic systematics model while assuming the input parameters given in Table 18. We find that the posteriors from KiDS-SBI accurately recover the true input parameters (see Table 21 for the inferred parameter values). As shown in Fig. 25, the anisotropic systematics posterior mostly only constrains $S_{8}$ and $A_{\mathrm{IA}}$. At the same time, it becomes apparent that the posterior distributions over $\omega_{\mathrm{b}}$, $n_{\mathrm{s}}$, $A_{\mathrm{bary}}$, and $h_{0}$ are mainly prior driven, namely, the posterior distribution is mostly flat throughout prior space. As the cosmic shear signal is largely degenerate to small changes in these parameters, this is expected.
Table 4: Main inferred cosmological and astrophysical parameters from mocks generated with the anisotropic systematics model.
@@ Appendix F Cosmology dependence of the cosmic shear two-point statistic covariance
To determine the origin of the observed cosmology dependence in the likelihood in Fig. 17, we investigate whether there is a theoretical justification for the uncertainty in the cosmic shear signal in KiDS-1000 to vary significantly with $S_{8}$. First, we assume that the likelihood of a full-sky shear angular power spectrum, $C(\ell)$, at a given cosmology, is Gaussian, specifically $\mathcal{L}\propto\mathrm{exp}(-\frac{1}{2}\chi^{2})$, where $\chi$ is the goodness of fit given by
@@
where $R^{(pq)}(\ell)\equiv C_{\mathrm{fid}}^{(pq)}(\ell)/\sqrt{C_{\mathrm{fid}}^{(pp)}(\ell)\,C_{\mathrm{fid}}^{(qq)}(\ell)}$. By calculating the distribution of $\mathrm{exp}(-\frac{1}{2}\chi_{a}^{2})$ over $\chi$ for different values of $\ell$ and $a$, we can assess the impact of rescaling the shear angular power spectrum directly on the Gaussian likelihood across different angular scales, $\ell$. To facilitate this, we can analytically compute the standard deviation of this distribution, $\sigma_{\chi_{a}}$, as follows
@@
Table 5: All inferred cosmological and astrophysical parameters varied within the anisotropic systematics and the standard isotropic systematics models from the KiDS-1000 gold sample.
@@
Appendix G Impact of the sampler on the Gaussian likelihood posteriors